# ClassDiffusion: More Aligned Personalization Tuning with Explicit Class Guidance

### v1

@@ 3.4 Semantic Preservation Loss
Recent personalization techniques are based on tuning and creating a link between reference images and a unique token denoted as <new1>. During the tuning process, the embeddings associated with this token are fine-tuned to align with the specific concept being targeted. When dealing with multiple concepts during training, a total of $N$ special tokens are utilized, with each one designated to represent a distinct concept. Our method is designed to reduce the semantic gap between the customized concepts and he center of the super-categories distribution in text space. Fig.  4 shows an overall pipeline of our method. By labeling the phrase containing a special token as $P_{SC}$ and the phrase containing only a class token as $P_{C}$, using $\mathbf{E}_{SC}$ and $\mathbf{E}_{C}$ to represent the embeddings outputs by the CLIP encoder of the personalized phrases and the class phrases respectively, the semantic preservation loss can be expressed by the following equation,
$\mathcal{L}_{cls}=\mathsf{\sum^{\mathit{N}}\sum^{\mathit{b}}\sum^{\mathit{L}}}D_{c}(\mathbf{E}_{SC},\mathbf{E}_{C})$ (10)
where $b$ represents for the batch size, $L$ denotes the length of embeddings for CLIP. $D_{c}$ implies the cosine distance. Therefore, the final optimization objective is:
$\mathcal{L}=\mathcal{L}_{recon}+\lambda\mathcal{L}_{cls}$ (11)
where $\lambda$ controls the weight of semantic preservation loss.
4 Experiments
Figure 6: Qualitative comparison between our method and Custom Diffusion(CD) in multiple concepts. Our method has better text similarity than custom diffusion.
4.1 Experiment Details
Implementation details Our method is built on Stable Diffusion V1.5, with a learning rate $10^{-6}$, and batch size 2 for fine-tuning. We use 500 optimization steps for single concept and 800 for multiple concepts, respectively. During inference, the guidance scale is set to 6.0, and inference steps are set to 100. All experiments are conducted on 2$\times$×RTX4090 GPUs. Our method uses $\sim$ 6 min for single concept generation and $\sim$ 11 min for multiple concept generation.
@@ 4.1 Experiment Details
Baselines We compare our method with state-of-the-art (SOTA) competitors including DreamBooth [64], Textual Inversion [20], Custom Diffusion [37], NeTI [1], SVDiff [29]. For DreamBooth, CustomDiffusion, and Textual Inversion, we used the diffusers [75] version of the implementation. For NeTI, we use its official implementation. Given SVDiff doesn’t have an official open-source repository. For SVDiff, we use the implementation of  [67]. All training parameters follow the recommendations of the official paper. To ensure the fairness of comparison, all of these baselines are built on Stable Diffusion V1.5.
@@
Datasets Following previous work [64, 29, 73], we conduct quantitative experiments on DreamBooth Dataset. It contains 30 objects including both live objects and non-live objects. Also, we use images from Textual Inversion Dataset and CustomConcept101 in qualitative experiments.
Figure 7: Qualitative result of a small story with given concepts.
@@
Evaluation metrics We assess our approach using three metrics: CLIP-I, CLIP-T, and DINO-I. CLIP-I calculates the visual similarity between the produced images and the target concept images by utilizing CLIP visual features. CLIP-T evaluates the similarity between text prompts and images. If one baseline contains the special token S*, it will be replaced with a prior class word. In the case of DINO-I, we evaluate the cosine similarity between the ViT-S/16 DINO [54] embeddings of the generated images and the concept images. Further, we note the impact of CLIP’s outdated performance on the fairness of the evaluation. Therefore we introduce BLIP2-T Score, which calculates the similarity between text features extracted from BLIP2’s Q-former and image features extracted from Vision Encoder as a score. Our approach involved utilizing the Transformer [80] implementation and the fine-tuned weights of BLIP2-IMT on CoCo [42], with ViT/L [15]. This new metric aims to offer a more equitable and efficient evaluation measure for future studies in this field. Empirical findings from various studies [39, 40, 81, 25, 84, 68] indicate that BLIP2 outperforms CLIP significantly in the assessment of text-image alignment.
@@ 4.2 Qualitative & Quantitative Experiments
Table 1: Quantitative Results on all metrics and Results of the User Study
@@
Single Concept DreamBooth [64] 0.249 0.855 0.700 0.295 95.4% 42.1%
Textual Inversion [20] 0.242 0.825 0.631 0.308 95.1% 75.0%
Custom Diffusion [37] 0.286 0.837 0.693 0.416 79.1% 40.0%
NeTI [1] 0.290 0.838 0.648 0.329 78.8% 70.0%
SVDiff [29] 0.293 0.834 0.606 0.418 56.6% 95.8%
[…]
Multiple Concept Custom Diffusion [37] 0.282 0.813 0.636 0.380 - -
[…]
Qualitative Experiments We compare our method with DreamBooth, Textual Inversion, NeTI, SVDiff, and Custom Diffusion on challenging prompts. The results depicted in Fig. 5 demonstrate the outcomes obtained from these prompts. Fig. 7 shows a story of a given dog and sunglasses. The experimental findings indicate a substantial superiority of our approach over other techniques regarding alignment with text prompts, without any decline in similarity to the specified concept. More qualitative results are shown in Appendix. G. Also, we conduct qualitative experiments with multiple concepts on the combinations <cat>, <sunglasses>; <bear>,<barn>; <dog>,<backpack>; Fig. 6 shows the results of the experiments. The experiments show that our method can be aligned with prompts better than custom diffusion in multi-concept generation.
@@
4.3 Additional Experiments
User Study We also performed user study to validate the effectiveness of our method. We used the same set of images generated in the Section 1 for user study, details of which are available in Appendix. F. The results of the user study are located in Tab. 1. The result of the user study shows that our method outperforms all methods in text similarity. Although our method does not outperform all methods in image similarity, given the high CLIP-I scores, our method still produces images that are highly consistent with the given concept.
@@ 5 Conclusions
In this work, we highlight the problem of weakened compositional ability due to individualized fine-tuning and provide an analysis of the causes of this problem from experimental observations and information-theoretic perspectives. We then introduce a new approach, termed ClassDiffusion, which mitigates the weakening of compositional ability by restoring the original semantic space. Finally, we present comprehensive experimental results showcasing the efficacy of ClassDiffusion and the fresh perspectives it offers on interconnected fields.
@@ Appendix A Visualization after Fine Tunning
Appendix B Fine-grained experiments
At the core of our approach, we want the semantics of personalized phrases to be closer to the category-centered words. In this section, we explore the effect of different category-centered words on the results for the same given concept. Fig. 11 shows the training results using different category-centered words. The results show that the use of different center words leads to significant differences in performance.
Figure 11: Visualization of generating “a <center word> sitting on the chair.”. This <center word> is used in both prompt and class token.
Appendix C Entropy reduction during the fine-tuning
@@ Appendix D Prompt used in the visualization of CLIP text sample space
In this section, we discuss how we visualized CLIP’s text sample space in Fig. 2. First, we use the adjectives generated which are described in Section. D as the initialize set $S$, and use the following pseudocode to get a 2D point set $T$:
@@
Appendix F Details of User Study
@@
Appendix G More Qualitative Result
@@ Appendix H Limitation
Appendix I Social Impact

### v3

@@ 3.4 Semantic Preservation Loss
Figure 5: The framework of ClassDiffusion. The personalization fine-tuning strategy is based on Custom Diffusion [39], which primarily fine-tunes the K and V parameters in the transformer block. Our semantic preservation loss (SPL) is calculated by measuring the cosine distance between text features extracted from the same text transformer (using EOS tokens as text features following CLIP) for phrases with personalized tokens and phrases with only super-class.
As analyzed above, the key challenge lies in preserving semantic information during fine-tuning to reduce the difficulty of sampling from the joint conditional distribution. To address this, we propose a novel loss function aimed at constraining semantic variation throughout the fine-tuning process.
Specifically, considering that there are $N$ special tokens, each representing a customized concept. During the process, the embeddings associated with these tokens are fine-tuned to align with the target concepts. Our loss function is designed to minimize the semantic distance between the phrase containing the special token (e.g., a photo of a V* dog) and the phrase containing only the class word (e.g., a photo of a dog). By labeling the training prompt as $P_{tp}$ (e.g., a photo of a V* dog), and the class prompt as $P_{cp}$ (e.g., a photo of a dog), we use a text encoder to get their embeddings $E_{tp}$ and $E_{cp}$ in Stable Diffusion’s semantic space. The semantic preservation loss (SPL) is calculated by the sum of the cosine distance of their text embeddings. Formally speaking, our proposed SPL can be expressed by the following equation:
$\mathcal{L}_{sp}=\mathsf{\sum^{\mathit{N}}\sum^{\mathit{B}}\sum^{\mathit{L}}}D_{c}\left(E_{SC},E_{C}\right)$ (10)
where $B$ represents the batch size, $L$ denotes the hidden dimensions of the text encoder, and $D_{c}$ implies the cosine distance. We can represent the final training objective as:
$\mathcal{L}=\mathcal{L}_{recon}+\lambda\mathcal{L}_{sp}$ (11)
The overview of our proposed model is shown in Fig. 5
4 Experiments
4.1 Experiment Details
Implementation details Our method is built on Stable Diffusion V1.5, with a learning rate $10^{-6}$, and batch size 2 for fine-tuning. We used 500 optimization steps for a single concept and 800 for multiple concepts, respectively. During inference, the guidance scale is set to 6.0 and the inference steps are set to 100. The semantical preservation loss weight is set to 1.0 during all experiments. All experiments are conducted on 2$\times$×RTX4090 GPUs. Our method uses $\sim$ 6 min for the generation of single concepts and $\sim$ 11 min for the generation of multiple concepts.
To better preserve the semantic space, we compute SPL between text embeddings embedded in the semantic space of the Stable Diffusion model. Therefore, we utilize the CLIP [61] text encoder from Stable Diffusion v1.5 [63], specifically clip-vit-large-patch14 [47], to extract the text embeddings of phrases. Following common practice, we use the End of Sequence (EOS) token to represent the semantics of embeddings.
@@ 4.1 Experiment Details
Baselines We compare our method with state-of-the-art (SOTA) competitors, including DreamBooth [66], Textual Inversion [20], Custom Diffusion [38], NeTI [1], SVDiff [29]. For DreamBooth, CustomDiffusion, and Textual Inversion, we used the diffusers [77] version of the implementation. For NeTI, we use its official implementation. Given that SVDiff does not have an official open-source repository. For SVDiff, we use the implementation of  [69]. All training parameters follow the recommendations of the official paper. To ensure fairness of comparison, all these baselines are built on Stable Diffusion V1.5.
@@
Datasets Following previous work [66, 29, 75], we conduct quantitative experiments on DreamBooth Dataset [66]. It contains 30 objects including both live objects and non-live objects. In addition, we used images from the Textual Inversion Dataset [20] and CustomConcept101 [38] in qualitative experiments.
@@
Evaluation metrics We assess our approach using three metrics: CLIP-I, CLIP-T, and DINO-I. CLIP-I calculates the visual similarity between the produced images and the target concept images by utilizing CLIP [61] visual features. CLIP-T evaluates the similarity between text prompts and images. If one baseline contains the special token S*, it will be replaced with a prior class word. In the case of DINO-I, we evaluate the cosine similarity between the ViT-S/16 DINO [56] embeddings of the generated images and the concept images. Further, we note the impact of CLIP’s outdated performance on the fairness of the evaluation. Therefore, we introduce the BLIP2-T Score, which calculates the similarity between text features extracted from BLIP2’s Q-former and image features extracted from Vision Encoder as a score. This metric is designed by calculating the similarity between image and text embeddings extracted by the BLIP2 model. Our approach involved utilizing the Transformer [82] implementation and the fine-tuned weights of BLIP2-IMT on CoCo [44], with ViT/L [15]. This new metric aims to offer a more equitable and efficient evaluation measure for future studies in this field. Empirical findings from various studies [40, 41, 83, 25, 86, 70] indicate that BLIP2 outperforms CLIP significantly in the assessment of text-image alignment.
@@ 4.2 Qualitative & Quantitative Experiments
Qualitative Experiments We compare our method with DreamBooth [66], Textual Inversion [20], NeTI [1], SVDiff [29], and Custom Diffusion [38] on challenging prompts. The results depicted in Fig. 6 demonstrate the outcomes obtained from these prompts. Fig. 2 shows a story of a given dog and sunglasses. The experimental findings indicate a substantial superiority of our approach over other techniques regarding alignment with text prompts, without any decline in similarity to the specified concept. More qualitative results are shown in Appendix. J. Also, we conduct qualitative experiments with multiple concepts on the combinations <cat>, <sunglasses>; <bear>,<barn>; <dog>,<backpack>; Fig. 7 shows the results of the experiments. The experiments show that our method can be aligned with prompts better than custom diffusion in multi-concept generation.
Table 1: Quantitative Results on all Metrics and Results of the User Study. The last two columns display the win rates of our method compared to other approaches in the user study, evaluated in terms of text similarity and image similarity. A win rate exceeding 50% (highlight by ) indicates that our method outperforms the compared methods on the corresponding metric, as judged from a human perspective.
@@
Single Concept DreamBooth [66] 0.249 0.855 0.700 0.295 0.559 95.4% 42.1%
Textual Inversion [20] 0.242 0.825 0.631 0.308 0.505 95.1% 75.0%
Custom Diffusion [38] 0.286 0.837 0.693 0.416 0.746 79.1% 40.0%
NeTI [1] 0.290 0.838 0.648 0.329 0.607 78.8% 70.0%
SVDiff [29] 0.293 0.834 0.606 0.418 0.835 56.6% 95.8%
Our 0.300 0.828 0.673 0.460 0.843 - -
Multiple Concepts Custom Diffusion [38] 0.282 0.813 0.636 0.380 - -
Our 0.320 0.821 0.604 0.477 - -
4.3 Additional Experiments
Figure 7: Qualitative comparison between our method and Custom Diffusion(CD) in multiple concepts. Our method has better text alignment than custom diffusion.
User Study We also performed user study to validate the effectiveness of our method. We used the same set of images generated in the Section 4.2 for user study, details of which are available in Appendix. I. The results of the user study are located in Tab. 1. The numbers in the table indicate at what percentage our method is considered by humans to be superior to the compared methods. The result of the user study shows that our method outperforms all methods in text similarity (> 50%). Although our method does not outperform all methods in image similarity, given the high CLIP-I scores, our method still produces images that are highly consistent with the given concept.
@@ 5 Conclusions
In this work, we highlight the problem of weakened compositional ability due to individualized fine-tuning and provide an analysis of the causes of this problem from experimental observations and information-theoretic perspectives. We discovered that this weakening effect is primarily attributed to the semantic shift of the customized concepts throughout the fine-tuning process. As the model undergoes fine-tuning, the representations of these concepts gradually drift away from their original meanings, leading to a misalignment with the intended semantics. This semantic drift complicates the model’s ability to accurately sample from the joint conditional distribution, ultimately hindering its performance in generating or understanding the intended outcomes based on the fine-tuned concepts. We then introduce a new approach, termed ClassDiffusion, which mitigates the weakening of compositional ability by restoring the original semantic space. Finally, we present comprehensive experimental results showcasing the efficacy of ClassDiffusion and the fresh perspectives it offers on interconnected fields.
6 Acknowledgement
This work was supported in part by the National Key R&D Program of China under Grant No. 2022YFC3310200, the National Natural Science Foundation of China (No.92470203, U23A20314), Beijing Natural Science Foundation (No. L242022 ), and the Fundamental Research Funds for the Central Universities (No. 2024XKRC082)
@@ Appendix A Visualization after Fine Tunning
Appendix B Entropy reduction during the fine-tuning
@@
Figure 11: Visualization of generating “a <center word> sitting on the chair.”. This <center word> is used in both prompt and class token. Experiments show that a fine-grained center word will benefit our proposed method.
Appendix C Fine-grained Experiments
At the core of our approach, we want the semantics of personalized phrases to be closer to the category-centered words. In this section, we explore the effect of different category-centered words on the results for the same given concept. Fig. 11 shows the training results using different category-centered words. The results show that the use of different center words leads to significant differences in performance, and a fine-grained center word benefits our method. Also, it indicates the importance of recovering the semantical space of the customized concept.
@@ Appendix D Prompt used in the visualization of CLIP text sample space
In this section, we discuss how we visualized CLIP’s text sample space in Fig. 3. First, we collect 70 phrases that combine adjectives with words representing the class of given concepts (e.g., "a happy dog" or "a cool dog"). Using the CLIP text encoder, we extract text embeddings for these phrases. To visualize the semantic space and intuitively track modifications within it, we use t-SNE to reduce these high-dimensional embedding vectors to 2D. An overview of this result are shown in Fig.3(a), meanwhile we show the real experimental results in Fig.12. Formally, we use the adjectives generated which are described in Section. D as the initialize set $S$, and use the following pseudocode to get a 2D point set $T$:
@@
Appendix F Multi-concepts Experiments
In this section, we demonstrate the ability of ClassDiffusion to generate multiple concepts, specifically 3 concepts in one model. Fig 13 shows the result of this experiment. The experiments demonstrate that ClassDiffusion generates high-quality results when combining multiple concepts, validating the effectiveness of our proposed methods.
Figure 13: Qualitative result of generating three concepts.
Appendix G More baseline comparison
To further evaluate the performance of our proposed method, we further introduce a new baseline Prospect [97] which aims at solving similar problems we observe. The result of the experiment is shown in Tab. 2. The quantitative results below show that our model achieves superior performance.
[…]
Prospect 0.294 0.815 0.588
ClassDiffusion(Ours) 0.300 0.828 0.673
Table 2: Quantitative results comparing with Prospect
Appendix H Quantitative ablation of SPL weight
In this section, we conduct a quantitative ablation experiment on the choose of SPL weight. Tab. 3 shows the result of the experiment result. This result shows that CLIP-T becomes higher (increase in the ability to follow prompts) and DINO-I decreases(decrease in the ability to customize concepts) with increasing SPL weight, which is consistent with our expectations. Meanwhile, we find that SPL is a loss function that is insensitive to weight. Combined with the qualitative observation in Fig. 8, we prefer to choose 1 as the SPL weight.
SPL weight CLIP-T $\uparrow$ DINO-I $\uparrow$
0.01 0.299 0.677
0.1 0.300 0.674
1 0.300 0.673
10 0.300 0.665
100 0.301 0.661
Table 3: Performance metrics for different SPL weights.
Appendix I Details of User Study
@@
Appendix J More Qualitative Result
@@
Appendix L Social Impact